sqrt(2x-5)-sqrt(x-3)=1
Answer by fz_shurtugal(3)

sqrt(2x-5)-sqrt(x-3)=1
Simplify this equation first, so basically you have sqrt(2x-5-x+3)=1
now simplify the radicand which would equall to sqrt(x-2),making the equation sqrt(x-2)=1.Now you want to get rid of the radical sign to solve for x, therefore you would square both sides of the equation, sqrt(x-2)^2=1^2,and then simplify, which would equall to, x-2=1. You now solve this equality which would be x=3.